OpenStudy (anonymous):

Log2(3x)+Log2(5)=Log2(30) what are the steps to take to solve this?

OpenStudy (anonymous):

use the log rule to combine the two logs on the left side of your equation...

OpenStudy (anonymous):

logb(x) + logb(y) = logb(x*y)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you got it from here or you want to continue??

OpenStudy (anonymous):

I think i got it. Is it Log2(3x•5)=Log2(30) Log2(15x)=Log2(30) 15x=30 x=2 ?

OpenStudy (anonymous):

nicely done! good job..